OpenStudy (anonymous):

The derivative of √x-1/x(x^1/3) is Answer is 1/2x^-1/2+4/3x^-7/3 Please explain how to get to this answer

OpenStudy (kc_kennylau):

\[\Large\sqrt x-\frac1x\sqrt[3]x\]

OpenStudy (kc_kennylau):

Is this what you mean?

OpenStudy (kc_kennylau):

I believe that it's \(\Large\displaystyle\sqrt x-\frac1{x\sqrt[3]x}\)

OpenStudy (kc_kennylau):

Convert everything to powers of x

OpenStudy (anonymous):

yes that's what I mean

OpenStudy (kc_kennylau):

\[\sqrt x=x^{\frac12}\]\[\sqrt[3]x=x^{\frac13}\]